Dan's father was forty years older when the son was 12 years old. How many years must be added to Dan's age so that he will be 8 years older than one-fourth of his father's age?
Answer by ankor@dixie-net.com(22740)

Dan's father was forty years older when the son was 12 years old. How many years must be added to Dan's age so that he will be 8 years older than one-fourth of his father's age?
:
"Dan's father was forty years older when the son was 12 years old."
:
Father's age = 12 + 40 = 52
Dan's age = 12
:
How many years must be added to Dan's age so that he will be 8 years older than one-fourth of his father's age?
:
Let x = no. of yr to be added
:
D + x = .25F + 8
Substitute the ages of F and D
12 + x = .25(52) + 8
12 + x = 13 + 8
x = 21 - 12
x = 9 yrs to be added
;
:
See if that is true in the statement:
" 9 years must be added to Dan's age so that he will be 8 years older than one-fourth of his father's age? "
12 + 9 = .25(52) + 8
21 = 21